A golf ball rolls off a horizontal cliff with an initial speed of 10.2 m/s. The ball falls a vertical distance of 15.3 m into a lake below. (a) How much time does the ball spend in the air? (b) What is the speed v of the ball just before it strikes the water?
Answer #1

Initial Speed, u = 10.2 m/s

height fallen through, H = 15.3 m

(A)

H = ut + \frac{1}{2}gt^2\\ or, 15.3 = 10.2*t + 0.5*9.8*t^2\\ or, t = \frac{-10.2\pm \sqrt{10.2^2-4*4.9*(-15.3)}}{2*4.9}\\ or,t = -3.1s, 1.0 s

Discarding negative value of t, time spend by the ball in the air is 1.0 s.

(b) v = u + gt

or, v = 10.2 + 9.8*1.0

or, v = 20.0 m/s